# Echo Chamber: RL Post-training Amplifies Behaviors Learned in Pretraining

### v1

@@ Abstract
Reinforcement learning (RL)-based fine-tuning has become a crucial step in post-training language models for advanced mathematical reasoning and coding. Following the success of frontier reasoning models, recent work has demonstrated that RL fine-tuning consistently improves performance, even in smaller-scale models; however, the underlying mechanisms driving these improvements are not well-understood. Understanding the effects of RL fine-tuning requires disentangling its interaction with pretraining data composition, hyperparameters, and model scale, but such problems are exacerbated by the lack of transparency regarding the training data used in many existing models. In this work, we present a systematic end-to-end study of RL fine-tuning for mathematical reasoning by training models entirely from scratch on different mixtures of fully open datasets. We investigate the effects of various RL fine-tuning algorithms (PPO, GRPO, and Expert Iteration) across models of different scales. Our study reveals that RL algorithms consistently converge towards a dominant output distribution, amplifying patterns in the pretraining data. We also find that models of different scales trained on the same data mixture will converge to distinct output distributions, suggesting that there are scale-dependent biases in model generalization. Moreover, we find that RL post-training on simpler questions can lead to performance gains on harder ones, indicating that certain reasoning capabilities generalize across tasks. Our findings show that small-scale proxies in controlled settings can elicit interesting insights regarding the role of RL in shaping language model behavior.111Our code is available at https://github.com/rosieyzh/openrlhf-pretrain.
@@ 2.2 Reinforcement Learning Fine-tuning
For the results presented in Section 3 we fine-tune using questions from the train split of GSM8K and study the performance and format of the generations of the models on the test split of GSM8K, both during and after fine-tuning. In Section 4 we take the models fine-tuned using questions from GSM8K and evaluate on the test set of MATH-500 and AIME 1983-2024. In Appendix H we also perform PPO on questions from the train split of MATH. For more details about the hyperparameters used, refer to Appendix C.
@@ 3 RL on Models Pretrained from Scratch with Different Mixtures
In this section, we present a summary of our results after applying reinforcement learning fine-tuning using problems from GSM8K on our models which were pretrained from scratch. With the exception of a few results in Section 3.3, we always include FineMath3+ and Algebraic-Stack in our pretraining mixtures, and vary quantities of TinyGSM, OpenMathInstruct1, and OpenMathInstruct2. Furthermore, unless otherwise specified, figures in this section correspond to our runs with PPO on models with 150M parameters; we conduct further analysis on models with 1B parameters in Section 3.4 and Appendix E, and comparisons with other RL algorithms and Expert Iteration are provided in Section 3.5 and Appendix F.
@@ 4 Transfer to other evaluation datasets
In Appendix G.1, we analyze these improvements from a qualitative lens by prompting GPT-4.5 Preview to classify the types of errors made by the base model for incorrect generations and later corrected following fine-tuning. In Appendix G.2 we present evaluation results on AIME for the same models and find little to no improvement on pass@1 and majority@64 performance for the AIME 2022–2024 benchmark across all pretrained models, but improvements are observed for pass@64 performance. In Appendix G.3 we provide examples of model generations on MATH-500 and AIME 2022-2024 before and after doing RL fine-tuning on GSM8K, where the base model was previously incorrect and the fine-tuned model provides a correct answer.
@@ Appendix A Related Works
Despite the large literature covering RL post-training, there is still a lack of understanding for the connection between the pretraining data and the effect it has on RL post-training optimization. To the best of our knowledge, we are the first to perform an extensive end-to-end study of the effect of pretraining data mixtures for mathematical reasoning in LLMs of different scales, and explore the difference between the common policy optimization algorithms. Havrilla et al. (2024) is the closest work to our own, studying the performance of PPO across scales both on base models and fine-tuned models. Pan et al. (2025) also explores the emergence of the “Aha” moment in base LLMs, trained for solving countdown and multiplication tasks. Finally, Gandhi et al. (2025) leverage continued pretraining on Llama models towards bringing their performance closer to the Qwen models, and show that this improvement correlates with the reasoning abilities of the initial model.
@@ F.2 Expert Iteration
Appendix G Further Transfer Learning Investigations
G.1 Qualitative Analysis on MATH-500 Generations
@@ G.2 AIME Evaluation
In Figure 22, we perform the same qualitative analysis on the generations for the AIME pass@64 evaluation as in Section G.1.
@@
G.3 Example Generations on MATH and AIME 2022-2024
@@
Appendix H RL Fine-tuning on MATH

### v2

@@ Abstract
Reinforcement learning (RL)-based fine-tuning has become a crucial step in post-training language models for advanced mathematical reasoning and coding. Following the success of frontier reasoning models, recent work has demonstrated that RL fine-tuning consistently improves performance, even in smaller-scale models; however, the underlying mechanisms driving these improvements are not well-understood. Understanding the effects of RL fine-tuning requires disentangling its interaction with pretraining data composition, hyperparameters, and model scale, but such problems are exacerbated by the lack of transparency regarding the training data used in many existing models. In this work, we present a systematic end-to-end study of RL fine-tuning for mathematical reasoning by training models entirely from scratch on different mixtures of fully open datasets. We investigate the effects of various RL fine-tuning algorithms (PPO, GRPO, and Expert Iteration) across models of different scales. Our study reveals that RL algorithms consistently converge towards a dominant output distribution, amplifying patterns in the pretraining data. We also find that models of different scales trained on the same data mixture will converge to distinct output distributions, suggesting that there are scale-dependent biases in model generalization. Moreover, we find that RL post-training on simpler questions can lead to performance gains on harder ones, indicating that certain reasoning capabilities generalize across tasks. Our findings show that small-scale proxies in controlled settings can elicit interesting insights regarding the role of RL in shaping language model behavior.111Our code is available at https://github.com/rosieyzh/openrlhf-pretrain. All pretrained base models can be found here, and intermediate checkpoints from RL fine-tuning for two 1B pretrained models can be found at the following links: TinyGSM + OMI1 + OMI2 and TinyGSM + OMI2.
@@ 2.2 Reinforcement Learning Fine-tuning
For the results presented in Section 3 we fine-tune using questions from the train split of GSM8K and study the performance and format of the generations of the models on the test split of GSM8K, both during and after fine-tuning. In Section 4 we take the models fine-tuned using questions from GSM8K and evaluate on the test set of MATH-500 and AIME 1983-2024. In Appendix I we also perform PPO on questions from the train split of MATH. For more details about the hyperparameters used, refer to Appendix C.
@@ 3 RL on Models Pretrained from Scratch with Different Mixtures
In this section, we present a summary of our results after applying reinforcement learning fine-tuning using problems from GSM8K on our models which were pretrained from scratch. With the exception of a few results in Section 3.3, we always include FineMath3+ and Algebraic-Stack in our pretraining mixtures, and vary quantities of TinyGSM, OpenMathInstruct1, and OpenMathInstruct2. Furthermore, unless otherwise specified, figures in this section correspond to our runs with PPO on models with 150M parameters; we conduct further analysis on models with 1B parameters in Section 3.4 and Appendix E, and comparisons with other RL algorithms and Expert Iteration are provided in Section 3.5 and Appendix F. Finally, we provide a brief theoretical justification of our results in Section 3.6.
@@ 3.1 RL converges to favour one distribution in the mixture
Finally, although we focus on accuracy and percentage metrics for our analysis here and henceforth in this section, we show that similar phenomena manifest even when tracking confidence-based metrics— such as the average probability of the TinyGSM and OpenMathInstruct1-style initial token formats— in Appendix G.
@@ 3.5 The effect of the RL algorithm
3.6 Supporting theory
We now provide some theoretical explanation for the results detailed above. We emphasize that the focus of this paper is not on theoretical analysis of reinforcement learning, and we simply reiterate known results that explain the findings of this work. Let be the space of inputs and be the space of responses. Let be a reward function, and let be our reference policy (before RL). Assume that our reference policy is in fact a mixture of different policies s.t. , for satisfying . For example, each can be a different solution format for math questions (code, text, etc.). We can frame the problem of reinforcement learning solved by e.g. PPO as maximizing the expected reward under KL-regularization222We note that our experimental results hold even without adding the KL-regularization term. We leave an analysis of this setting to future work.:
Then, the maximizer would correspond to:
Namely, we reweight the original mixture of policies corresponding to the rewards from each policy in the original mixture. This is consistent with our experiments, which show that RL mostly converges to the strategy which maximizes the reward.
@@ 4 Transfer to other evaluation datasets
In Appendix H.1, we analyze these improvements from a qualitative lens by prompting GPT-4.5 Preview to classify the types of errors made by the base model for incorrect generations and later corrected following fine-tuning. In Appendix H.2 we present evaluation results on AIME for the same models and find little to no improvement on pass@1 and majority@64 performance for the AIME 2022–2024 benchmark across all pretrained models, but improvements are observed for pass@64 performance. In Appendix H.3 we provide examples of model generations on MATH-500 and AIME 2022-2024 before and after doing RL fine-tuning on GSM8K, where the base model was previously incorrect and the fine-tuned model provides a correct answer.
@@ Appendix A Related Works
Despite the large literature covering RL post-training, there is still a lack of understanding for the connection between the pretraining data and the effect it has on RL post-training optimization. To the best of our knowledge, we are the first to perform an extensive end-to-end study of the effect of pretraining data mixtures for mathematical reasoning in LLMs of different scales, and explore the difference between the common policy optimization algorithms. A theoretical explanation for the diversity collapse brought by RLVR is presented in Wu et al. (2025), who argue that RLVR is inherently limited to the support of the base model. Havrilla et al. (2024) is the closest work to our own, studying the performance of PPO across scales both on base models and fine-tuned models. Pan et al. (2025) also explores the emergence of the “Aha” moment in base LLMs, trained for solving countdown and multiplication tasks. Finally, Gandhi et al. (2025) leverage continued pretraining on Llama models towards bringing their performance closer to the Qwen models, and show that this improvement correlates with the reasoning abilities of the initial model.
@@ F.2 Expert Iteration
Appendix G Confidence-Based Metrics
Our results in Section 3 highlight how different pretraining data mixtures influence both the stylistic distribution and accuracy of model outputs. We now show that these preferences also manifest in confidence-based metrics.
During RL fine-tuning, we track the average probability of outputs beginning with def simple_math_problem() and Let’s solve this problem using Python code. <llm-code> on the GSM8K test set. As detailed in Appendix B, these token prefixes are characteristic of TinyGSM and OMI1-style generations, respectively. (We exclude OMI2 from this analysis due to the lack of a consistent initial token pattern.) As shown in Figure 22, the average probabilities closely follow the trends in output proportions presented in Figures 2, 4(a), and 4(b), albeit with a smoother trajectory. Additionally, the narrowing error bars over the course of training suggest further stability.
Overall, we found that the average generation probabilities increase throughout training—even after the output format has largely stabilized—indicating that the model’s confidence continues to grow within the dominant output distribution.
Appendix H Further Transfer Learning Investigations
H.1 Qualitative Analysis on MATH-500 Generations
@@ H.2 AIME Evaluation
In Figure 23, we perform the same qualitative analysis on the generations for the AIME pass@64 evaluation as in Section H.1.
@@
H.3 Example Generations on MATH and AIME 2022-2024
@@
Appendix I RL Fine-tuning on MATH